# AGE-Net: Spectral–Spatial Fusion and Anatomical Graph Reasoning with Evidential Ordinal Regression for Knee Osteoarthritis Grading

### v1

@@ Appendix A Additional implementation notes
A.1 Training tricks and stability
We adopt mixed precision training to improve throughput, and EMA to stabilize evaluation. Layer-wise learning rate decouples backbone and head updates: the backbone uses a smaller learning rate to reduce catastrophic forgetting of pretrained features, while the head learns task-specific mapping. We disable ranking loss under Mixup because labels become interpolated and the ordering constraint becomes ill-defined in soft label space.
@@ A.2 Test-time augmentation
We apply flip-based TTA by averaging predictions of the original and horizontally flipped inputs. This can reduce variance and improve robustness, especially for noisy radiographs.
Appendix B Planned extended ablations
w/o SSF: remove frequency modulation and spatial gating to isolate spectral contribution.
@@ Appendix B Planned extended ablations
w/o COE: replace evidential head with heteroscedastic regression head and compare uncertainty quality.
k sensitivity: vary k in AGR (e.g., 5/9/13) to study neighborhood size effects.
Token grid size: vary pooling grid (e.g., 1010, 1414, 1616).
Appendix C Statistical significance (planned)
We will report paired significance tests across seeds (e.g., paired -tests on per-seed QWK) and effect sizes when comparing full AGE-Net with major ablations and the ConvNeXt baseline.

### v2

@@ Appendix A Additional implementation notes
A.1 Training details and stability
We used automatic mixed precision to improve training throughput and reduce memory consumption. To stabilize evaluation and reduce epoch-to-epoch variance, we maintained an exponential moving average (EMA) of model parameters and reported results using the EMA weights at test time. Because the backbone was initialized from pretrained ConvNeXt weights, we adopted layer-wise learning-rate decoupling: the backbone was optimized with a smaller learning rate to mitigate catastrophic forgetting of transferable representations, whereas the newly introduced modules and prediction head used a larger learning rate to learn task-specific mappings. Unless otherwise specified, all ablations followed the same optimization schedule to ensure fair comparison.
We also found it important to handle interactions between augmentation and ordinal constraints. In particular, we disabled the ranking loss when Mixup augmentation was active because Mixup produces interpolated (soft) labels, making the strict inequality ill-defined and potentially introducing inconsistent supervision in the ordinal pair set. In such cases, the model was trained using the evidential regression objective alone.
@@ A.2 Test-time augmentation
At inference, we applied a lightweight flip-based test-time augmentation (TTA) by averaging predictions from the original image and its horizontally flipped counterpart. This procedure can reduce prediction variance and improve robustness to minor pose and acquisition differences without increasing training complexity. For the evidential head, we averaged the predicted severity scores and used the corresponding averaged score for final grade rounding; uncertainty-related parameters were handled consistently with the score aggregation used in our implementation.
Appendix B Extended ablations and statistical analysis
B.1 Planned extended ablations
To further characterize the contribution of each design choice and assess robustness of hyperparameter settings, we plan the following additional ablations:
w/o SSF: remove frequency modulation and spatial gating to isolate the contribution of spectral–spatial enhancement.
@@ B.1 Planned extended ablations
w/o COE: replace the evidential NIG head with a conventional heteroscedastic regression head and compare predictive uncertainty behavior.
sensitivity: vary in AGR (e.g., 5/9/13) to study the effect of neighborhood size on non-local reasoning.
Token grid size: vary the pooling grid resolution (e.g., , , ) to evaluate the trade-off between token granularity and computational cost.
B.2 Planned statistical analysis
To quantify the reliability of observed improvements, we plan to report paired comparisons across random seeds when evaluating AGE-Net against major ablations and the ConvNeXt baseline. Specifically, we will apply paired statistical tests on per-seed summary metrics (e.g., paired -tests on QWK) and report corresponding effect sizes. Where appropriate, we will additionally compute confidence intervals (e.g., via bootstrap resampling) to complement -values and better convey the magnitude and uncertainty of performance differences.